OpenStudy (anonymous):

A reaction has an equilibrium constant of 8.6×103 at 298 K. At 677K , the equilibrium constant is 0.59. Find ΔHrxn for the reaction._______ kJ

OpenStudy (anonymous):

So this is solved by Vant Hoff's equation: \[\ln \left( \frac{ (Kp)1 }{ (Kp)2 } \right)=-\frac{ DeltaH }{ R }\left( \frac{ 1 }{ T1 }-\frac{ 1 }{ T2 } \right)\] which would get us dH=42,429.24J/mol or 42.429 KJ/mol

OpenStudy (aaronq):

^good job! however, you should let the user try to solve the question for themselves.

OpenStudy (anonymous):

@aaronq ok jaja